A particular fruit's weights are normally distributed, with a mean of 747 grams and a standard deviation of 39 grams.

The heaviest 13% of fruits weigh more than how many grams?

Give your answer to the nearest gram.
Answer #1

Solution:-

Given that,

mean = = 747

standard deviation = = 39

Using standard normal table,

P(Z > z) = 13%

= 1 - P(Z < z) = 0.13  

= P(Z < z) = 1 - 0.13

= P(Z < z ) = 0.87

= P(Z < 1.13 ) = 0.87  

z = 1.13

Using z-score formula,

x = z * +

x = 1.13 * 39 + 747

x = 791.07

x = 791 gram